# Simulating Chemistry on Bosonic Quantum Devices

### v1

@@ 4.1 Elementary chemical reaction dynamics
Figure 8: The correspondence between the bosonic cQED platform device and the chemical systems via a double-well isomorphism. The bosonic cQED device (top left corner, adapted from Ref. 48) can be used to perform analog simulation of chemistry (top right), according to the mapping between the Hamiltonian parameter space of the Kerr-Cat Hamiltonian (middle left) and a generic chemical double-well (middle right). The bottom right figure showcases state functions of the device Hamiltonian plotted in a phase space representation (adapted from Ref. 48), which can be thought of as analogous to the states associated with a chemical double-well potential.
@@
Bosonic quantum platforms such as the cQED-based Kerr-cat devices offer a general strategy for exploring the chemical dynamics of reactions that can be described by a double-well potential energy surface. 48 The cQED platforms allow tuning dissipation for accurate modeling of chemical reactivity in dissipative environments. Given that chemical dynamics is inherently quantum-mechanical in nature, simulating chemical dynamics on bosonic devices provides a novel avenue for exploring chemical reactivity, complementary to traditional approaches based on spectroscopic methods and numerical simulations.
@@
The adenine-thymine proton transfer reaction 82 is an example of a biologically significant chemical reaction that can be modeled in terms of a double-well free energy profile, 83 as shown in Figure 8. Simulating the chemical dynamics of the adenine-thymine proton transfer covers a wide range of double-well free energy profiles since it changes significantly during the cell life cycle. Indeed, the system can be modeled as a chemical double-well potential ($\hat{U}=k4\hat{x}^{4}-k_{2}\hat{x}^{2}+k_{1}\hat{x}$), where each of the parameters $k_{4},k_{2},k_{1}$ are related to the energy landscape from reactants to products, as shown in Figure 8.
@@
The quantum dynamics of the bosonic Kerr-cat cQED device described here have been extensively studied. 84, 23, 85, 86 This experimental platform can be thought of as a double-well analog, as shown in Figure 8, with kinetic behavior that depends on the drives of the system as well as the inherent dissipation of the physical experiment. 48, 87. The parameters of the device can be continuously varied, something that is difficult to achieve in an experimental chemical setting. Furthermore, given the mapping between the device and chemical double-well Hamiltonians, this cQED platform offers the possibility to study a wide range of chemical systems. Thus, the flexibility and the presence of inherent quantum effects in cQED devices can extend and complement the toolset available for studying chemical reaction dynamics, while alleviating the difficulty and labor-intensive efforts of physical chemistry experiments.
@@ 4.2 Photochemical reaction dynamics and conical intersections
Figure 9: Two-state, two-dimensional potential energy surface that describes photoinduced isomerization of retinal. Reprinted from Ref. 88.
@@
Figure 10: A proposed cQED circuit for the simulation of a conical intersection between two molecular energy surfaces such as in Figure 9. The artificial chemical system is realizable by a superconducting circuit, comprised of three Josephson junctions (indicated by red boxes) 89 describing the electronic properties of the molecules, whereas the green charged node describes the electronic state of the system. The quantronium circuit is coupled to a set of three microwave superconducting cavities, encoding the vibrational degrees of freedom of the molecules. The system dynamics are driven by pulses that modulate the excitation of the electronic and vibrational modes. The coupling between the quantronium and cavities can act as a surrogate for the vibronic coupling of the chemical system around the conical intersection.
@@
Conical intersections are formed when electronic states corresponding to different molecular geometries become energetically degenerate and are critical for simulating photochemical reactions. 90, 91 Recently, an experimental demonstration of an engineered conical intersection in a hybrid cQED processor for a minimal model of a two-dimensional linear vibronic coupling has shown promise in the use of bosonic quantum devices as tunable simulators of chemical reactivity. 92 For the model considered, experimental measurements captured the delicate interplay between coherent evolution and dissipation that is ubiquitous in chemical reactivity. By engineering a cQED device with five simultaneous microwave drives and tunable dissipation and performing Wigner state tomography, branching of the wavepacket was observed when dissipative coupling induced electronic excitation corresponding to jumps between potential energy surfaces. 92
@@
The success in engineering the conical intersection with a cavity-transmon device motivates further studies on exploring nonadiabatic chemical dynamics through conical intersections with a bosonic cQED machine. For an initial exploration, a model system that describes the ultrafast photoisomerization of retinal molecules might be most suitable. 88 This two-surface, two-dimensional model system describes the $S_{0}$ and $S_{1}$ electronic surfaces, each containing two coordinates, namely, the isomerization dihedral angle reaction coordinate and a polyene stretching coordinate, as shown in Figure 9. The coupling between the two surfaces is linear in the stretching coordinate, drawing proximity with the model system in Ref. 92. However, different from that system, the retinal reactive coordinate is modeled with a cosine function for describing the cis-trans isomerization and is therefore highly anharmonic. 88 It is possible to represent the retinal model system with a Hamiltonian of individual transmons, that encode the electronic states, and a set of microwave cavities that encode vibrational degrees of freedom. Thus, we envision the retinal potential energy surface to be effectively represented with a transmon-based device such as the one shown in Figure 10. 89 Hence, we demonstrate how this platform can facilitate the investigation of chemical dynamics around conical intersections that harness the capabilities of the bosonic cQED machines.
@@ 4.3 Molecular graph problems
Figure 11: Molecular substructure search of compounds that contain the query structure. Apomorphine only admits one such substructure, whereas others contain at least two. The first and second substructure’s atoms are in blue and yellow, respectively, where the overlapping atoms are in pink. It has been observed that the biosynthesis of morphine in opium puppy, 93 starting from (R)-reticuline through three other intermediates, preserves the two substructures (not necessarily their conformations) that are similar to the query.
@@
The advent of big data has led to the mass digitizing of information, including molecular data relevant to chemistry and biosciences. Naturally represented by graphs, molecules (and thus chemical reactions) can be stored electronically as massive databases, such as PubChem 94, SciFinder, RCSB Protein Data Bank 95, just to name a few. It should be of no surprise that solving graph problems provides quantitative insights into our knowledge of the chemical properties of molecules, proteins, and macromolecules. Many graph problems, however, are known to be computationally challenging on larger scales. For instance, molecular substructure search and retrosynthetic analysis rely on efficiently finding common subgraphs, which is closely related to the subgraph isomorphism problem, as shown in Figure 11. However, solving the subgraph isomorphism problem is impractical on classical computers beyond a certain size since it is a nondeterministic polynomial-time (NP) problem. 96
@@
One promising direction to tackle this challenge is to harness the power of photonic quantum computing, which has been demonstrated to gain a quantum advantage in the Gaussian boson sampling (GBS) problem. 97 GBS is a quantum computing approach that utilizes photonic Gaussian states passing through linear interferometers. Despite its simplicity, GBS has the potential for solving important molecular graph problems and molecular docking, efficiently simulating phenomena intractable for classical computers.98 Unlike other types of quantum computers which struggle with high degrees of entanglement, photonic circuits are capable of creating large quantum systems with millions of entangled modes via time-domain multiplexing techniques where the number of optical devices used is independent of the number of qumodes. 97 Whereas graph-like structures are ubiquitous in quantum information and for generating or manipulating entanglement, 99, 100 encoding graph characteristics into physical modes allows for efficient analog simulation of complex high-dimensional properties in these bosonic quantum systems. Exploring the power of GBS beyond known graph problems, i.e., towards optimization 101 also represents fundamental directions, as well as exploring well-known heuristics for graph problems in the quantum setting. Understanding or mitigating noise 102 due to photon loss or other factors in these sampling procedures is also critical for realizing advantages.
@@
To apply the state-of-the-art bosonic cQED quantum platform to solve important molecular graph problems with GBS, a reuse-and-repurpose approach can be taken for efficiently implementing large (a few tens of optical modes) GBS problems with small, experimentally-accessible cQED devices with only a handful of cavity modes. 103 This scheme was tested with the problem of binding a thiol-containing aryl sulfonamide ligand to the tumor necrosis factor-$\alpha$ converting enzyme receptor, showing that the approach facilitates efficient simulation of large (50-100 modes) GBS problems with cQED devices that contain only a handful of modes. 103
@@ 4.4 Electronic structure calculations
Simulation of a molecular electronic structure on bosonic quantum devices must include a key step of mapping this fermionic problem to a bosonic one. Traditional quantum computers also tackle a similar problem, where the electronic structure must be mapped to a system of qubits. In fact, a Fock space description of qubits reveals that qubits are parafermions: particles with hybrid boson-fermion quantum statistics. 104 In the case of qumodes, electronic structure mapping can be achieved by mapping the excitations between molecular orbitals to the excitations between the levels in qumodes. 105 Specifically, a molecular system of $N$ electrons and $M$ spin-orbitals can be mapped to a system of $N$ qumodes with $M-N+1$ levels. As a specific example, we show how electronic states of molecular orbitals, also known as Slater determinants, can be mapped to Fock states of two qumodes for the dihydrogen molecule in Figure 12. This allows for writing the electronic structure Hamiltonian in the Fock basis of the qumodes, which in turn allows the computation of the molecular electronic energy by a set of Fock number measurements on the cQED device. 22
@@ 4.5 Qudit quantum phase estimation using qumodes
A qudit is a generalization of the qubit in quantum computing. While a qubit can exist in states $|0\rangle$ and $|1\rangle$, a qudit can exist in states $\{|0\rangle,|1\rangle,\cdots,|d-1\rangle\}$, where $d$ is the dimension of the qudit. So, a qubit is essentially a special case of a qudit where $d=2$. Many of the qubit quantum algorithms such as the Quantum Fourier Transform (QFT) and Quantum Phase Estimation (QPE) (as shown in Figure 13) can be generalized to their qudit versions. 106 Extending these qubit algorithms to qudits reduces the required number of qudits and qudit gates.107 As the QPE has been applied extensively in quantum chemistry simulations such as electronic structure calculations, the generalization to qudits can benefit these simulations and extend them to more complex chemical systems.
While qudits and qumodes represent different aspects of quantum systems—qudits deal with discrete quantum states while qumodes deal with CV quantum states—one could encode qudit states into qumodes. This is done by encoding these states into different coherent states of the electromagnetic field. 108 Specifically, we could use $d$ different coherent states, each corresponding to one of the qudit states. A 2-qutrit QPE has been demonstrated with a procedure developed for designing the circuit for qudit QPE. 109, 106 Encoding qudit states into qumodes allows us to implement more qudits with higher dimensions and thus expand the overall computational space. Implementing the qudit quantum phase estimation algorithm using qumodes can be used for obtaining the eigenvalues of a given Hamiltonian such as the potential energy surfaces of molecules.
@@ 5 Concluding remarks
Fault-tolerant quantum computing (FTQC), based on quantum error correction (QEC), is the only reliable and credible path toward scalable large-scale general-purpose applications. 110, 111 However, FTQC requires substantial qubit and gate resources. 111 Open-loop quantum control methods such as dynamical decoupling (DD) 112 and its many variants are thus attractive in this regard, as has been demonstrated in recent experiments using transmons. 113, 114, 115, 116, 117, 118 Under appropriate conditions, integrating DD with a QEC scheme can lead to a reduced accuracy threshold for FTQC. 119 In practice, a DD-protected gate (that is, a gate executed following a DD pulse) can achieve greater accuracy than an unprotected gate. A critical challenge remains: to validate DD methods on cQED platforms by demonstrating that the effective noise level for a DD-protected gate is indeed lower than for an unprotected gate. Successfully meeting this challenge would mark a substantial stride towards an accelerated realization of FTQC.
@@
We acknowledge support from the NSF-CCI Center for Quantum Dynamics on Modular Quantum Devices (CQD-MQD) under grant number CHE-2124511.

### v3

@@ 4.1 Elementary chemical reaction dynamics
Figure 8: The correspondence between the bosonic cQED platform device and the chemical systems via a double-well isomorphism. The bosonic cQED device (top left corner, adapted from Ref. 57) can be used to perform analog simulation of chemistry (top right), according to the mapping between the Hamiltonian parameter space of the Kerr-Cat Hamiltonian (middle left) and a generic chemical double-well (middle right). The bottom right figure showcases state functions of the device Hamiltonian plotted in a phase space representation (adapted from Ref. 57), which can be thought of as analogous to the states associated with a chemical double-well potential.
@@
In this section, we give an outlook regarding the asymmetric Kerr-cat as it relates to chemistry, originally from Ref. 93 and Ref. 94. Bosonic quantum platforms such as the cQED-based Kerr-cat devices offer a general strategy for exploring the chemical dynamics of reactions that can be described by a double-well potential energy surface. 57 The cQED platforms allow tuning dissipation for accurate modeling of chemical reactivity in dissipative environments. Given that chemical dynamics is inherently quantum-mechanical in nature, simulating chemical dynamics on bosonic devices provides a novel avenue for exploring chemical reactivity, complementary to traditional approaches based on spectroscopic methods and numerical simulations.
@@
The adenine-thymine proton transfer reaction 95 is an example of a biologically significant chemical reaction that can be modeled in terms of a double-well free energy profile, 96 as shown in Figure 8. Simulating the chemical dynamics of the adenine-thymine proton transfer covers a wide range of double-well free energy profiles since it changes significantly during the cell life cycle. Indeed, the system can be modeled as a chemical double-well potential ($\hat{U}=k4\hat{x}^{4}-k_{2}\hat{x}^{2}+k_{1}\hat{x}$), where each of the parameters $k_{4},k_{2},k_{1}$ are related to the energy landscape from reactants to products, as shown in Figure 8.
@@
The quantum dynamics of the bosonic Kerr-cat cQED device described here have been extensively studied. 97, 24, 98, 99 This experimental platform can be thought of as a double-well analog, as shown in Figure 8, with kinetic behavior that depends on the drives of the system as well as the inherent dissipation of the physical experiment. 57, 100. The parameters of the device can be continuously tuned, something that is difficult to achieve in an experimental chemical setting due to the correlated chemical properties of different functional groups and rich chemical microenvironment. Furthermore, given the mapping between the device and chemical double-well Hamiltonians, this cQED platform offers the possibility to study a wide range of chemical systems, and perform high-throughput chemical kinetic studies without the limitations of synthetic chemistry or the numerical limitations of exact open quantum dynamics methods.93, 94 This is of greater relevance for chemical problems for which multiple molecular degrees of freedom are relevant for reactivity for which coupled cQED devices could address the numerical limitations of high-dimensionality open quantum dynamics methods. Thus, the flexibility and the presence of inherent quantum effects in cQED devices can extend and complement the toolset available for studying chemical reaction dynamics, while alleviating the difficulty and labor-intensive efforts of physical chemistry experiments.
@@ 4.2 Photochemical reaction dynamics and conical intersections
Figure 9: Two-state, two-dimensional potential energy surface that describes photoinduced isomerization of retinal. Reprinted from Ref. 101.
@@
Figure 10: A proposed cQED circuit for the simulation of a conical intersection between two molecular energy surfaces such as in Figure 9. The artificial chemical system is realizable by a superconducting circuit, comprised of three Josephson junctions (indicated by red boxes) 102 describing the electronic properties of the molecules, whereas the green charged node describes the electronic state of the system. The quantronium circuit is coupled to a set of three microwave superconducting cavities, encoding the vibrational degrees of freedom of the molecules. The system dynamics are driven by pulses that modulate the excitation of the electronic and vibrational modes. The coupling between the quantronium and cavities can act as a surrogate for the vibronic coupling of the chemical system around the conical intersection.
@@
Conical intersections are formed when electronic states corresponding to different molecular geometries become energetically degenerate and are critical for simulating photochemical reactions. 103, 104 Recently, an experimental demonstration of an engineered conical intersection in a hybrid cQED processor for a minimal model of a two-dimensional linear vibronic coupling has shown promise in the use of bosonic quantum devices as tunable simulators of chemical reactivity. 105 For the model considered, experimental measurements captured the delicate interplay between coherent evolution and dissipation that is ubiquitous in chemical reactivity. By engineering a cQED device with five simultaneous microwave drives and tunable dissipation and performing Wigner state tomography, branching of the wavepacket was observed when dissipative coupling induced electronic excitation corresponding to jumps between potential energy surfaces. 105
@@
The success in engineering the conical intersection with a cavity-transmon device motivates further studies on exploring nonadiabatic chemical dynamics through conical intersections with a bosonic cQED machine. For an initial exploration, a model system that describes the ultrafast photoisomerization of retinal molecules might be most suitable. 101 This two-surface, two-dimensional model system describes the $S_{0}$ and $S_{1}$ electronic surfaces, each containing two coordinates, namely, the isomerization dihedral angle reaction coordinate and a polyene stretching coordinate, as shown in Figure 9. The coupling between the two surfaces is linear in the stretching coordinate, drawing proximity with the model system in Ref. 105. However, different from that system, the retinal reactive coordinate is modeled with a cosine function for describing the cis-trans isomerization and is therefore highly anharmonic. 101 It is possible to represent the retinal model system with a Hamiltonian of individual transmons, that encode the electronic states, and a set of microwave cavities that encode vibrational degrees of freedom. The anharmonicity makes this an example of a numerically challenging problem, and thus obtaining an exact open quantum dynamics description both at short and long times could give insight into the photophysics of the problem. Thus, a transmon-based device could provide a bridge between state-of-the-art numerical methods and existing photochemical experiments, by explicit inclusion of term-by-term quantum effects that inherently incorporate the entanglement of multiple potential energy surfaces and other contributing environmental degrees of freedom. Thus, we envision the retinal potential energy surface to be effectively represented with a transmon-based device such as the one shown in Figure 10, 102 and serve as a proof of concept for analogous problems that may be challenging to exactly simulate on a classical device due to the dimensionality of the problem.
Hence, we demonstrate how this platform can facilitate the investigation of chemical dynamics around conical intersections that harness the capabilities of the bosonic cQED machines.
@@ 4.3 Molecular graph problems
Figure 11: Molecular substructure search of compounds that contain the query structure. Apomorphine only admits one such substructure, whereas others contain at least two. The first and second substructure’s atoms are in blue and yellow, respectively, where the overlapping atoms are in pink. It has been observed that the biosynthesis of morphine in opium poppy, 106 starting from (R)-reticuline through three other intermediates, preserves the two substructures (not necessarily their conformations) that are similar to the query.
@@
The advent of big data has led to the mass digitizing of information, including molecular data relevant to chemistry and biosciences. Naturally represented by graphs, molecules (and thus chemical reactions) can be stored electronically as massive databases, such as PubChem 107, SciFinder, RCSB Protein Data Bank 108, just to name a few. It should be of no surprise that solving graph problems provides quantitative insights into our knowledge of the chemical properties of molecules, proteins, and macromolecules. Many graph problems, however, are known to be computationally challenging on larger scales. For instance, molecular substructure search and retrosynthetic analysis rely on efficiently finding common subgraphs, which is closely related to the subgraph isomorphism problem, as shown in Figure 11. However, solving the subgraph isomorphism problem, a nondeterministic polynomial-time (NP) problem 109, generally relies on efficient heuristics to speed up existing algorithms, reduction of subgraph isomorphism tests, or relaxing isomorphism conditions 110 that is inherently hard on classical computers 111, especially for large graphs.
One promising direction to tackle this challenge is to harness the power of photonic quantum computing, which has been demonstrated to gain a quantum advantage in the Gaussian boson sampling (GBS) problem. 112
@@
Pioneered over a decade ago, 113 boson sampling (BS) was proposed as a non-universal quantum computation model to compute the permanent of matrices which is a # P-hard problem. One photonic implementation of BS is GBS, which samples from squeezed input Gaussian states, a non-classical resource, after being passed through linear interferometers. In doing so, the GBS protocol approximates the hafnians of matrices, related to the number of perfect matchings of graphs,114, 41 rather than permanents obtained from BS. Hence, GBS has the potential for solving important molecular graph problems and molecular docking, efficiently simulating phenomena intractable for classical computers.115 Unlike other types of quantum computers which struggle with high degrees of entanglement, photonic circuits are capable of creating large quantum systems with millions of entangled modes via time-domain multiplexing techniques where the number of optical devices used is independent of the number of qumodes. 112 Whereas graph-like structures are ubiquitous in quantum information and for generating or manipulating entanglement, 116, 117 encoding graph characteristics into physical modes allows for efficient analog simulation of complex high-dimensional properties in these bosonic quantum systems. Exploring the power of GBS beyond known graph problems, i.e., towards optimization 118 also represents fundamental directions, as well as exploring well-known heuristics for graph problems in the quantum setting. Understanding or mitigating noise 119 due to photon loss or other factors in these sampling procedures is also critical for realizing advantages.
@@
To apply the state-of-the-art bosonic cQED quantum platform to solve important molecular graph problems with GBS, a reuse-and-repurpose approach can be taken for efficiently implementing large (a few tens of optical modes) GBS problems with small, experimentally-accessible cQED devices with only a handful of cavity modes. 120 This scheme was tested with the problem of binding a thiol-containing aryl sulfonamide ligand to the tumor necrosis factor-$\alpha$ converting enzyme receptor, showing that the approach facilitates efficient simulation of large (50-100 modes) GBS problems with cQED devices that contain only a handful of modes. 120
The realistic implementation of boson sampling (and thus GBS) only provides approximate sampling, due to imperfections such as decoherence, noise, and photon losses. Thus, the complexity of GBS algorithms should also be compared to their approximate classical counterparts. In light of recently proposed classical sampling algorithms 121 for graph problems, the development of GBS algorithms is even more relevant for the exploration of the power of GBS in general, as well as to encourage a further expansion of the scope of quantum-inspired algorithms.
@@ 4.4 Electronic structure calculations
Simulation of a molecular electronic structure on bosonic quantum devices must include a key step of mapping this fermionic problem to a bosonic one. Traditional quantum computers also tackle a similar problem, where the electronic structure must be mapped to a system of qubits, which can be understood as spin-$1/2$ systems or hardcore bosons. In fact, a Fock space description of qubits reveals that qubits are parafermions: particles with hybrid boson-fermion quantum statistics. 122 In the case of qumodes, electronic structure mapping can be achieved by mapping the excitations between molecular orbitals to the excitations between the levels in qumodes. 123 Specifically, a molecular system of $N$ electrons and $M$ spin-orbitals can be mapped to a system of $N$ qumodes with $M-N+1$ levels. 124 As a specific example, we show how electronic states of molecular orbitals, also known as Slater determinants, can be mapped to Fock states of two qumodes for the dihydrogen molecule in Figure 12. The state mapping can then be used to map the bilinear fermionic operator $E_{q}^{p}=f_{p}^{\dagger}f_{q}$ in terms of qumode Fock space projection operators of the form |$\ket{n_{1},\cdots,n_{N}}\bra{m_{1},\cdots,m_{N}}$ |, where $f_{p}^{\dagger}$ and $f_{q}$ are fermionic creation and annihilation operators, respectively. 125, 124 The operator mapping is exact and thus preserves the antisymmetry of fermionic algebra. The operator mapping only includes a subset of all projection operators possible in the corresponding qumode Fock space, thus the number of terms scales polynomially with system size. 124 This allows for writing the electronic structure Hamiltonian in the Fock basis of the qumodes, which in turn allows the computation of the molecular electronic energy by a set of Fock number measurements on the cQED device. 23 A bosonic ansatz has also recently been explored for molecular electronic structure combined with fermion to qubit mapping.126
@@ 4.5 Qudit quantum phase estimation using qumodes
Qumodes, with their expanded Hilbert space dimensionality, are natural candidates for representing qudits. A qudit is a generalization of the qubit in quantum computing. While a qubit can exist in states $|0\rangle$ and $|1\rangle$, a qudit can exist in states $\{|0\rangle,|1\rangle,\cdots,|d-1\rangle\}$, where $d$ is the dimension of the qudit. So, a qubit is essentially a special case of a qudit where $d=2$. Many of the qubit quantum algorithms such as the Quantum Fourier Transform (QFT) and Quantum Phase Estimation (QPE) (as shown in Figure 13) can be generalized to their qudit versions. 127 Extending these qubit algorithms to qudits reduces the required number of qudits and qudit gates.128 As the QPE has been applied extensively in quantum chemistry simulations such as electronic structure calculations, the generalization to qudits can benefit these simulations and extend them to more complex chemical systems. Generalizing the QPE to the qudit representation provides a notable improvement in terms of reducing the number of required qudits and decreasing the error rate exponentially as the qudit dimension increases 128. Applications of qudit-based QPE beyond simulating complex chemical systems is a growing area of research. For example, a method to solve the linear system using a qutrit-based QPE algorithm has been proposed 129.
While qudits and qumodes represent different aspects of quantum systems—qudits deal with discrete quantum states while qumodes deal with CV quantum states—one could encode qudit states into qumodes. This is done by encoding these states into different coherent states of the electromagnetic field. 130 Specifically, we could use $d$ different coherent states, each corresponding to one of the qudit states. A 2-qutrit QPE has been demonstrated with a procedure developed for designing the circuit for qudit QPE. 131, 127 Encoding qudit states into qumodes allows us to implement more qudits with higher dimensions and thus expand the overall computational space. Implementing the qudit quantum phase estimation algorithm using qumodes can be used for obtaining the eigenvalues of a given Hamiltonian such as the potential energy surfaces of molecules.
@@ 5 Concluding remarks
Despite the advantages of the described bosonic cQED platforms, some of the outstanding challenges include increasing the number of coupled qumodes and the number of states that can be resolved within each qumode.23 The latter point is important for efficiently and accurately encoding chemical problems, such as those involved in dynamics and electronic structure problems, while the former is necessary for answering higher-dimensional chemical questions. One approach has been to utilize universal bosonic circuits to encode algorithmic operations. For example, there has been recent progress in developing universal bosonic circuits based on the cQED platform, that can in principle represent any unitary of arbitrary dimension, including gates based on multiple qubits. This can open the door for quantum algorithms that can use the multiple levels of one oscillator rather than using multiple qubits. 132, 133, 61, 60 Thus, we can encode and study chemical problems within a compact bosonic framework, aiding to the development of enhanced, high-quantum-volume bosonic devices with scalable error-tolerance.
Fault-tolerant quantum computing (FTQC), based on quantum error correction (QEC), is the only reliable and credible path toward scalable large-scale general-purpose applications. 134, 135 However, FTQC requires substantial qubit and gate resources. 135 Open-loop quantum control methods such as dynamical decoupling (DD) 136 and its many variants are thus attractive in this regard, as has been demonstrated in recent experiments using transmons. 137, 138, 139, 140, 141, 142 Under appropriate conditions, integrating DD with a QEC scheme can lead to a reduced accuracy threshold for FTQC. 143 In practice, a DD-protected gate (that is, a gate executed following a DD pulse) can achieve greater accuracy than an unprotected gate. A critical challenge remains: to validate DD methods on cQED platforms by demonstrating that the effective noise level for a DD-protected gate is indeed lower than for an unprotected gate. Successfully meeting this challenge would mark a substantial stride towards an accelerated realization of FTQC.
Parallel to the progress in hardware and algorithms, software libraries 144, 145, 146, 147 have been developed that can numerically emulate bosonic quantum devices on classical computers. However, these software capabilities are still at a nascent stage compared to their qubit counterparts, and still have room for growth. Publicizing software libraries alongside comprehensive documentation with tutorials will contribute to the expansion of the bosonic quantum computing community, expediting the availability and usefulness of this quantum platform for applications within and beyond chemistry.
@@
We acknowledge support from the NSF-CCI Center for Quantum Dynamics on Modular Quantum Devices (CQD-MQD) under grant number CHE-2124511. Heidi Hendrickson acknowledges computational resources provided in part by the MERCURY consortium under NSF grants CHE-1229354, CHE-1662030, and CHE-2018427.